OpenStudy (anonymous):

Part 1: Write the equation of the line that passes through the points (–1, –4) and (2, 5) in point-slope form. Part 2: Using complete sentences, explain whether or not it matters which point is used in the final answer. Also explain why you chose the point you did.

OpenStudy (anonymous):

first you need to find the slope... slope(m) = (y2 - y1) / (x2 - x1) (-1,-4) x1 = -1, and y1 = -4 (2,5) x2 = 2, and y2 = 5 can you find the slope by subbing in the info into the formula ?

OpenStudy (anonymous):

Yes. @kelliegirl33

OpenStudy (anonymous):

slope(m) = (y2 - y1) / (x2 - x1) m = (5 - (-4) / (2-(-1) m = (5 + 4) / (2 + 1) m = 9/3 m = 3 Once you find the slope, you will then need to use point slope formula to find the equation. y - y1 = m(x - x1) use slope(m) = 3 and either of the points : (-1,-4) x1 = -1, y1 = -4) (2,5) x1 = 2, y1 = 5 y - (-4) = 3(x - (-1) y - 5 = 3(x - 2) y + 4 = 3(x + 1) y - 5 = 3x - 6 y + 4 = 3x + 3 y = 3x - 6 + 5 y = 3x + 3 - 4 y = 3x - 1 y = 3x - 1 As you can see, either point you use will give you the same answer.